OpenStudy (anonymous):

a ladder leans against a vertical wall with the bottom of the ladder 8 ft from the wall on a horizontal floor. at that time the bottom end of the ladder is being pulled away at the rate of 3ft/s and the top of the ladder slips down the wall at a rate of 4 ft/s. how long is the ladder?

OpenStudy (anonymous):

look a related rate ladder problem. 

OpenStudy (anonymous):

\[z^2=x^2+y^2\] \[0=2xx'=2yy'\] \[0=xx'+yy'\] put \[x=8,x'=3,y'=4\] and solve for y then get the length of the ladder by pythagoras btw the left hand side is 0 because the length of the ladder is constant